# Amandine Brunetto  Sascha Hornauer  Fabien Moutarde Centre for Robotics

### v1

@@ 6 Discussion
A primary limitation of our method is its scene dependency, requiring NeRAF to be trained separately for each scene. Future work should explore advancements in generalizable NeRF to develop a unified, generalizable approach for synthesizing audio-visual scenes. Additionally, due to the absence of a suitable real-world dataset for our task, our method has not been tested on real data. Future research will address this by evaluating the method in real-world scenarios. The current approach is also limited to static scenes with a single sound source at a time. Addressing the challenge of learning implicit representations for audio-visual scenes with multiple and dynamic sound sources would be a significant advancement. Finally, our work relies on RIRs, which, while effective at capturing scene acoustics, are not easy to obtain. It would be beneficial to evaluate the method using more common or ambient sounds.
Societal Impact.
@@ Conclusion.
We introduce NeRAF, a cross-modal method that learns both neural radiance and acoustic fields. By conditioning the acoustic field with scene features from a 3D voxel grid containing color and density information from NeRF, NeRAF enables realistic audio auralization, spatialization and novel view synthesis. Our results demonstrate its promise: substantial improvement over other methods, enhanced novel view synthesis in complex scenes, increased data efficiency and easier access to audio-visual synthesis.
@@
We would like to thank Simon de Moreau for his help and insights. This work was supported by the French Agence Nationale de la Recherche (ANR), under grant ANR22-CE94-0003.
Supplementary Material
Appendix A Architectures
@@
NeRAF works with any NeRF method without modifications. It only requires a radiance field that the grid sampler can query using voxel center coordinates and viewing directions. Our method only relies on NeRF to obtain a voxel-grid representation of the scene. In this paper, we choose Nerfacto from Nerfstudio [53]. For more details on its architecture we invite readers to refer to Nerfstudio website and paper.
@@
We train a ResNet3D-50 to embed the grid into 2,048 features that best condition the acoustic field. At inference, only these features, and not the complete grid, are required to perform RIR synthesis. Note that we employ average pooling at the end of the ResNet, adjusting according to the grid size to consistently yield 2,048 features. NAcF consists of 2 MLP blocks. The first MLP block takes as input the multimodal fusion of grid features, encoded positions, and directions obtained through vector concatenation. It comprises 5 layers, each followed by a Leaky ReLU activation with a slope of 0.1, and outputs a 512 intermediate representation of the acoustic field. The second MLP block takes this intermediate representation as input and predicts the $F$ frequencies corresponding to the time query, aiming to learn the HRTF. It consists of one separate MLP layer per microphone channel. The final activation layer is a tanh function scaled between $[-10,10]$ to fit the STFT log-magnitude range.
Appendix B Implementation Details
@@
We implement our method using PyTorch framework [37]. We optimize NAcF using Adam optimizer [26] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ and $\epsilon=10^{-15}$. The initial learning rate is $10^{-4}$. It decreases exponentially to reach $10^{-8}$. For NeRF, we keep default Nerfacto parameters.
@@
For the first $2,000$ iterations, we only train the NeRF part. It allows the grid to be filled and updated several times using batches of $4,096$ voxel-centers. After, both NeRF and NAcF are train jointly. We use batch sizes of $4,096$ for NeRF and $2,048$ for NAcF. NeRAF is trained for 500k iterations but most runs reach their peak performance before, depending of the room size. We train our method on a single RTX 4090 GPU.
@@
Similar to previous works, we resample all RIRs from 44,100 Hz to 22,050 Hz. We compute STFT with 512 FFT bins, a Hann window of size 512 and hop length of 128. We obtain log-magnitude STFT using $\log(|\text{STFT}|+10^{-3})$. Like [30], we cut STFT to a maximum length depending of the scene. If the STFT is shorter than the maximum length we pad it with its minimal value. We inverse the transform of the predicted STFT using PyTorch Griffin-Lim algorithm running on GPU. We compute metrics against ground-truth waveforms.
@@
We give 3D microphone and source positions to NAcF. They go through multi-scale positional encoding with $N=10$ frequencies and maximum frequency exponent of $e=8$. Microphone rotation is scaled between $[0,2\pi]$ Then, we apply positional encoding with $N=4$ and $e=8$. The time queries are normalized within the range $[-1,1]$ and go through positional encoding with $N=10$ and $e=8$.
For the learning objective defined in Section 4.5, we select $\lambda_{A}=10^{-3}$, $\lambda_{SC}=10^{-1}$ and $\lambda_{SL}=1$.
To evaluate the quality of synthesized RIR waveforms, $w$, we use T60, C50 and EDT errors obtained as follows:
@@
$\text{EDT}(w,w_{\text{GT}})=|\text{EDT}(w)-\text{EDT}(w_{\text{GT}})|$ (13)
@@
To evaluate NeRF performances we choose metrics commonly used such as PSNR, SSIM and LPIPS.
Appendix C Scene Views Generation
@@
SoundSpaces 1.0 comes with a code to generate RGB and depth observations at microphone positions and orientations. They corresponds to $128\times 128$ resolution images with a $60^{\circ}$ field-of-view (FOV) sampled every 0.5 m with orientations $\in[0^{\circ},90^{\circ},180^{\circ},270^{\circ}]$. In NeRAF, we decided to not rely on them for NeRF training and instead generate our own observations using Habitat Sim. This is motivated by real scenarios constraints. Views of the space may be obtained via a video recorded in the room. It is unlikely that someone will stop every 0.5 m and turn around to obtain images at each orientations. Moreover, those kind of views are not optimal for NeRF training leading, in the worst case, to the failure of less sophisticated methods, and, in the best case, to the use of more data. We also wanted to decorrelate microphone poses and camera poses for greater freedom in the usage of the method.
@@
With Habitat Sim we rendered higher resolutions images with size $512\times 512$ and with a $90^{\circ}$ FOV. The larger FOV ensures more overlap between views. We randomly select positions at the edge of the room, orienting the camera to face its center with a random offset. For small room, medium and large room we use respectively 45, 75 and 150 training images. 50 test poses are randomly sampled in the room. Figure S1 presents examples of the positions obtained.
@@
Figure S1: Locations of visual observations for 3 SoundSpaces scenes. Blue dots corresponds to training views and green crosses to evaluation views.
Appendix D Number of Parameters, Storage and Inference Speed
@@
We provide in Table S1 the numbers of parameters of the NeRAF model and evaluate its inference storage requirements and speed. The speed was averaged over the 6 rooms on a single RTX 4090. The storage is the same for every rooms as they all use a 2,048-features representation of the scene regardless of its size. Note that at inference, the grid is not necessary and the extracted features are sufficient allowing the method to be faster and more compact.
@@
Table S1: Model number of parameters along with storage and speed requirements at inference.
Parameters (Million) Storage (MB) Speed (ms)
71.85 97.70 7.04
Appendix E Grid Visualization
@@
We display examples of sectional views of grids obtained using the grid sampler in Figure S2. They are then encoded into features using ResNet, which condition our neural acoustic fields.
@@
Figure S2: Sectional views of $128^{3}$ resolution grids obtained using the grid sampler.
Appendix F Additional Results Visualization
@@
We provide examples of log-magnitude STFTs predicted with NeRAF in Figure S3
@@
We render binaural RIRs at various distances from the sound source and perform auralization by convolving an anechoic sound with these RIRs. This demonstrates one of the primary applications of NeRAF. The chosen sound is open source and sourced from AVAD-VR [54]. In Figure S4, we illustrate that NeRAF is distance-aware: as the microphone approaches the sound source, the amplitude increases, and it decreases as the microphone moves away. Additionally, when the microphone is positioned to the right or left of the source, the audio channels’ amplitudes reflect this spatial positioning.
@@
We provide additional qualitative results on the effect of joint learning neural acoustic and radiance fields in Figure S5.
Figure S3: Examples of predicted log-magnitudes of binaural RIR STFTs compared to corresponding ground truth (GT). L and R respectively stand for left and right.
Figure S4: Examples of sounds generated with NeRAF on frl apartment 2. Our method successfully synthesizes spatialized and distance-aware sounds.
Figure S5: Cross-modal learning additional visualization.
Appendix G License Information
Asset License
Soundspaces [6] Creative Commons Attribution 4.0 International License
Habitat-Sim [45] MIT License
Nerfstudio [53] Apache License 2.0
Replica [50] Replica license

### v2

@@ 5 Discussion
A primary limitation of our method is its scene dependency, requiring NeRAF to be trained separately for each scene. Future work should explore advancements in generalizable NeRF to develop a unified, generalizable approach for synthesizing audio-visual scenes. The current approach is also limited to static scenes. Addressing the challenge of learning implicit representations for audio-visual scenes with dynamic sound sources would be a significant advancement.
@@ Conclusion.
We introduce NeRAF, a cross-modal method that learns both neural radiance and acoustic fields. By conditioning the acoustic field with scene features from a 3D grid containing geometric and appearance information, NeRAF enables realistic audio auralization, spatialization and novel view synthesis. Our results demonstrate improvements over other methods, enhanced novel view synthesis in complex scenes and increased data efficiency.
@@
We would like to thank Simon de Moreau for his help and insights. This work was supported by the French Agence Nationale de la Recherche (ANR), under grant ANR22-CE94-0003 and was performed using HPC resources from GENCI–IDRIS (Grant 2024-103789).
Appendix A Field parametrization
SoundSpaces.
SoundSpaces dataset lacks the complete acoustic field parameterization described in Equation 5. Microphone rotates only around the up-axis, $\theta$, and the source is omni-directional. Thus, NAcF can be expressed as $(\textbf{X}_{m},\textbf{d}_{m}=\theta,\textbf{X}_{s},t)\rightarrow(\text{RIR}_{\text{ch}1},\text{RIR}_{\text{ch}2})(f,t)$.
RAF.
RAF dataset contains only monaural RIRs. Consequently, we replace the two heads presented in Figure 4 by a single head. The microphone is omni-directional but the source is directional. We can therefore express NAcF as $(\textbf{X}_{m},\textbf{d}_{s}=\theta,\textbf{X}_{s},t)\rightarrow\text{RIR}_{\text{mono}}(f,t)$.
Appendix B Architectures
@@
NeRAF works with any NeRF method without modifications. It only requires a radiance field that the grid sampler can query using voxel center coordinates and viewing directions. Our method only relies on NeRF to fill a voxel-grid representation of the scene. In this paper, we choose Nerfacto from Nerfstudio (Tancik et al., 2023). For more details on its architecture we invite readers to refer to Nerfstudio website and paper.
@@
We train a ResNet3D-50 to embed the grid into features that best condition the acoustic field. For smaller scenes (office 4, room 2, apartment 2 and RAF dataset), we rely on 1,024 features. For larger scenes (frl apartment 2, 4, apartment 2) we use 2,048 features. At inference, only these features, and not the complete grid, are required to perform RIR synthesis. Note that we employ average pooling at the end of the ResNet, adjusting according to the grid size to consistently yield the desired number of features.
@@
NAcF consists of 2 MLP blocks. The first MLP block takes as input the multimodal fusion of grid features, encoded positions, and directions obtained through vector concatenation. It comprises 5 layers, each followed by a Leaky ReLU activation with a slope of 0.1, and outputs a 512 intermediate representation of the acoustic field. The second MLP block takes this intermediate representation as input and predicts the $F$ frequencies corresponding to the time query, aiming to learn the microphone specifics (e.g., HRTF for binaural microphones in SoundSpaces). It consists of one separate MLP head per microphone channel. The final activation layer is a tanh function scaled between $[-10,10]$ to fit the STFT log-magnitude range.
Appendix C Implementation Details
Hyperparameters.
@@ Appendix C Implementation Details
We implement our method using PyTorch framework (Paszke et al., 2019). We optimize NAcF using Adam optimizer (Kingma & Ba, 2014) with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ and $\epsilon=10^{-15}$. The initial learning rate is $10^{-4}$. It decreases exponentially to reach $10^{-8}$. For NeRF, just as AV-NeRF we keep default Nerfacto parameters.
@@
For the first 2k iterations, we only train the NeRF part. It allows the grid to be filled and updated several times using batches of $4,096$ voxel-centers. After, both NeRF and NAcF are train jointly. We use batch sizes of $4,096$ for NeRF and $2,048$ for NAcF. Note that NeRAF is trained by shuffling all STFT time slices in the train set. In NeRAF++, we require the full STFT to apply the energy decay loss so we train NeRAF by querying all the time bins necessaries to form the STFT. In that case we use a batch size of 34 STFTs. NeRAF is trained for 500k iterations but most runs reach their peak performance before, depending of the room size. We train our method on a single RTX 4090 GPU.
Learning objective details.
For the learning objective defined in Section 3.5, we empirically select $\lambda_{A}=10^{-3}$, $\lambda_{SC}=10^{-1}$ and $\lambda_{SL}=1$.
To train NeRAF++, we add an additional energy decay loss term that reinforces the energy attenuation tendency of predicted RIRs. Note that unlike INRAS++ and NACF, we do not use multi-resolution STFT and only apply the loss to the predicted STFT. The audio learning objective becomes:
$\mathcal{L}_{A}=\lambda_{SC}\mathcal{L}_{SC}+\lambda_{SL}\mathcal{L}_{SL}+\lambda_{ED}\mathcal{L}_{ED}.\$ (11)
with $\lambda_{SC}=3$, $\lambda_{SL}=1.5$ and $\lambda_{ED}=5$.
We follow Liang et al. (2023b) formula. Given a magnitude STFT $\mathbf{M}$, we first compute its energy in each time window by calculating the magnitude’s square and aggregating it along the frequency dimension:
[…]
where $\mathbf{M}^{\prime}\in\mathbb{R}_{+}^{D\times 2}$ represents the energy in each time window. We then sum $\mathbf{M}^{\prime}$ along the time dimension, aggregating the energy from the current step $d$ until the end $D$ for each time step $d\in[1,D]$ to capture the overall energy decay trend:
[…]
where $\mathbf{M}^{\prime\prime}\in\mathbb{R}_{+}^{D\times 2}$. Finally, we measure the L1 distance between the ground-truth energy trend and the predicted one in the log space:
$\mathcal{L}_{\mathrm{ED}}=||\log_{10}\mathbf{M}^{\prime\prime}_{g}-\log_{10}\mathbf{M}^{\prime\prime}_{p}||_{1}\enspace.$ (14)
SoundSpaces.
@@
To align with previous works on SoundSpaces, we resample all RIRs from 44,100 Hz to 22,050 Hz. We compute STFT with 512 FFT bins, a Hann window of size 512 and hop length of 128. We obtain log-magnitude STFT using $\log(|\text{STFT}|+10^{-3})$. Like Luo et al. (2022), we cut STFT to a maximum length depending of the scene. If the STFT is shorter than the maximum length we pad it with its minimal value.
RAF.
According to RAF (Chen et al., 2024), we cut audio to $0.32$ s and keep RIRs at 48 kHz. We compute STFT using 1,024 FFT bins, a Hann window of size 512 and hop length of 256.
Queries Encoding.
@@
We give the 3D position of microphone and source to NAcF. They are normalized using minimum and maximum values of the poses in the training set. We add a 1 m margin. Then, they undergo multi-scale positional encoding with $N=10$ frequencies and maximum frequency exponent of $e=8$. We obtain direction vectors from the microphone rotation and apply spherical harmonic encoding with 4 levels. The time queries are normalized within the range $[0,1]$ and go through positional encoding with $N=10$ and $e=8$.
Evaluation Metrics.
We inverse the transform of the predicted STFT using PyTorch Griffin-Lim algorithm running on GPU. We compute metrics against ground-truth waveforms.
To evaluate the quality of synthesized RIR waveforms $w$, we use T60, C50 and EDT errors obtained as follows:
@@
$\text{EDT}(w,w_{\text{GT}})=|\text{EDT}(w)-\text{EDT}(w_{\text{GT}})|.$ (17)
For RAF dataset, we follow previous works and also assess the quality of the time-frequency representation using the STFT error. To ensure fair comparison with methods directly predicting RIRs in the time domain, we perform the metric on the STFT obtained from the RIR in time domain. It means that we first go from the STFT predicted by NeRAF to a RIR waveform through Griffin-Lim algorithm and then compute its STFT before applying the metric. The STFT error is expressed as the absolute error between the predicted and the ground-truth log-magnitude STFTs:
STFT error ($\text{STFT error }(\mathbf{M},\mathbf{M}_{GT})=|\log(\mathbf{M})-\log(\mathbf{M}_{GT})|$ (18)
@@
To evaluate NeRF performances we choose the commonly used metrics such as PSNR, SSIM and LPIPS.
Appendix D Scene Views
SoundSpaces.
@@ Appendix D Scene Views
SoundSpaces 1.0 provides a code to generate RGB and depth observations at microphone positions and orientations. They corresponds to $128\times 128$ resolution images with a $60^{\circ}$ field-of-view (FOV) sampled every 0.5 m with orientations $\theta\in[0^{\circ},90^{\circ},180^{\circ},270^{\circ}]$. In NeRAF, we decided to not rely on them for NeRF training and instead generate our own observations using Habitat Sim. This is motivated by real scenarios constraints. Views of the space may be obtained via a video recorded in the room. It is unlikely that someone will stop every 0.5 m and turn around to obtain images at each orientations. Moreover, those kind of views are not optimal for NeRF training leading, in the worst case, to the failure of less sophisticated methods, and, in the best case, to the use of more data. We also wanted to decorrelate microphone poses and camera poses for greater freedom in the usage of the method.
@@
With Habitat Sim we rendered higher resolutions images with size $512\times 512$ and with a $90^{\circ}$ FOV. The larger FOV ensures more overlap between views. We randomly select positions at the edge of the room, orienting the camera to face its center with a random offset. For small room, medium and large room we use respectively 45, 75 and 150 training images. 50 test poses are randomly sampled in the room. Figure A1 presents examples of the positions obtained.
@@
Figure A1: Locations of visual observations for 3 SoundSpaces scenes. Blue dots corresponds to training views and green crosses to evaluation views. The exact poses coordinates are given in the supplementary material.
RAF.
Real Acoustic Fields dataset captures high-quality and dense room impulse response data paired with multi-view images using VR-NeRF ”Eyeful Tower” camera rig (Xu et al., 2023). It contains 22 cameras at different heights and pointing at different directions. We chose cameras 20 to 25 for training NeRAF’s vision part. This choice ensures that we have sufficient overlapping views. The images have a 0.7 megapixels with resolution (i.e., 684 × 1024). For more details on those data we invite readers to refer to VR-NeRF paper. To train NeRAF, as we want to limit the number of images necessaries, we randomly sub-select one third of those images. We let Nerfstudio select 90% of them for training and 10% for evaluation.
Figure A2: Evaluation of NeRAF on RAF with a decreasing number of training data.
Appendix E Few-shot on RAF.
RAF is a real-world dataset that contains a more realistic number of training data compared to SoundSpaces. We show in Figure A2 that NeRAF still outperforms INRAS, NAF and AV-NeRF with 25% less training data making it more data efficient.
Appendix F Number of Parameters, Storage and Inference Speed
@@
We provide in Table A1 the numbers of parameters of the NeRAF audio model and evaluate its inference storage requirements and speed. We present results when the 1,024 and 2,048 features are extracted from the grid. For RAF dataset, all results presented in the paper are obtained using 1,024 features while for SoundSpaces half the scenes use 1,024 (office 4, room 2, apartment 2) and the other half 2,048 (frl 2, frl 4, apartment 1). The speed was averaged over the RAF dataset on a single RTX 4090. Note that at inference, the grid is not necessary and the extracted features are sufficient allowing the method to be faster and more compact. As our inference speed time is two orders of magnitude smaller than the length of the generated audio, NeRAF can be used for real-time applications.
@@
Table A1: Audio model number of parameters along with storage and speed requirements at inference.
Grid Features Parameters (Million) Storage (MB) Speed (ms)
1024 55.49 77.93 2.33
2048 71.89 97.84 2.64
Appendix G Societal Impact
Appendix H Grid Visualization
@@
We display examples of sectional views of grids obtained using the grid sampler in Figure A3. They are then encoded into features using ResNet, which condition our neural acoustic fields.
@@
Figure A3: Sectional views of $128^{3}$ grids obtained using the grid sampler.
Figure A4: Examples of predicted log-magnitudes of binaural RIR STFTs compared to corresponding ground truth (GT). L and R respectively stand for left and right.
Appendix I Additional Results Visualization
@@
We provide examples of log-magnitude STFTs predicted with NeRAF in Figure A4.
@@
We render binaural RIRs at various distances from the sound source and perform auralization by convolving an anechoic sound with these RIRs. This demonstrates one of the primary applications of NeRAF. In Figure A5, we illustrate that NeRAF is distance-aware: as the microphone approaches the sound source, the amplitude increases, and it decreases as the microphone moves away. Additionally, when the microphone is positioned to the right or left of the source, the audio channels’ amplitudes reflect this spatial positioning.
Video.
We provide video examples of NeRAF generation in supplementary material. The chosen audios are open source and sourced from AVAD-VR (Thery & Katz, 2019).
@@
We provide additional qualitative results on the effect of joint learning neural acoustic and radiance fields in Figure A6.
Figure A5: Examples of sounds generated with NeRAF on frl apartment 2. Our method successfully synthesizes spatialized and distance-aware sounds.
Figure A6: Additional visualization of cross-modal learning benefits on the visual modality.
Appendix J RWAVS Dataset
Although we are aware AV-NeRF releases and evaluates its method on the RWAVS dataset, we chose to not present those results in the main paper. This is motivated by the difference between the tasks. NeRAF focus on learning RIRs, allowing the auralization of any given audio at inference. However, the RWAVS dataset does not contain RIRs but recordings of audio signals (e.g., music) played in a given environment. On this dataset, the network is therefore supposed to implicitly learn RIRs from them. To the best of our knowledge, it was never demonstrated that the representation learned can be used for the auralization of new audios.
Despite this, we present NeRAF performances on this dataset in Table A2. Similar to AV-NeRF, we adapt NeRAF to query frequencies instead of time and learn STFT masks. The same inverse modifications were done by AV-NeRF to evaluate on Sound Spaces (please see their supplementary material for more details). For every indoor scenes we extract 4,096 features from the grid and 1,024 for outdoors. This difference can be explained by the complexity of indoor scenes (e.g., many objects) compared to outdoors in this dataset. Because our loss combination is tailored for explicit RIR learning, we switch to MSE between magnitude STFTs.
Overall, NeRAF outperforms previous works tailored for RIR predictions and outperforms AV-NeRF in office, outdoors and slightly on apartment (ENV). We believe NeRAF is less performant on House because of the unusually large number of small objects in the scene. It would need a very high-resolution grid to capture them.
RWAVS contains very complex environments on which we can show that NeRAF cross-modal training benefits vision. Results are presented in Table A3. This demonstrates that learning both acoustics and radiance fields is interesting on both real and simulated data.
Table A2: Comparison with state-of-the-art methods on RWAVS dataset. Bold is best, underlined is second best.
Methods Office House Apartment Outdoors Overall
MAG ENV MAG ENV MAG ENV MAG ENV MAG ENV
INRAS 1.405 0.141 3.511 0.182 3.421 0.201 1.502 0.130 2.460 0.164
NAF 1.244 0.137 3.259 0.178 3.345 0.193 1.284 0.121 2.283 0.157
ViGAS 1 1.049 0.132 2.502 0.161 2.600 0.187 1.169 0.121 1.830 0.150
AV-NeRF 0.930 0.129 2.009 0.155 2.230 0.184 0.845 0.111 1.504 0.145
NeRAF 0.876 0.125 2.312 0.156 2.366 0.184 0.829 0.110 1.596 0.144
  • 1
    1Chen et al., Novel-view Acoustic Synthesis, CVPR 2023
Table A3: Quantitative results on RWAVS large complex scenes with sparse training data.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Training Images 1/6 1/4 1/6 1/4 1/6 1/4
House
Nerfacto 18.20 18.70 0.674 0.708 0.339 0.270
NeRAF 18.43 18.85 0.685 0.710 0.324 0.264
$\Delta\text{NeRAF}$ +1.26% +0.81% +1.65% +0.25% -4.31% -2.14%
Apartment
Nerfacto 19.14 19.64 0.746 0.776 0.251 0.211
NeRAF 19.24 19.68 0.748 0.778 0.249 0.210
$\Delta\text{NeRAF}$ +0.51% +0.23% +0.23% +0.23% -0.96% -0.34%